### v2

@@ 1 Introduction
  • We present PoseCrafter, a one-shot pose-guided personalized video generation framework with various inference techniques to improve sample quality.
@@ 2.1 Controllable Image Generation
Recently, because of the capacity to produce high-fidelity samples, diffusion models are inherently utilized in controllable image generation [47, 14, 8]. Text-prompt-based methods [30, 29, 32] often struggle with precise control, limited by text-image pair quality and complexity of lengthy inputs. Attention control [14, 8] and latent editing [27, 5] can somewhat enhance alignment of text-to-image generation. Personalized methods [31, 17] capture the common concept across a few images and reproduce that concept in generated images. Furthermore, some research [45, 25] introduces adapters that input various conditions like skeleton images to steer semantics and spatial structures. Our work centers on controllable video generation with pre-trained adapters and latent editing.
@@ 2.3 Personalization Video Generation
Text-to-video models [15, 35, 22] create a specific subject video through text description. Image-to-video approaches [1, 10, 26] animate a still image to produce a dynamic variation about it. To harness the vast array of personalized image models in the community, certain techniques [13] transform these into video generation models by incorporating temporal modeling modules pre-trained on a large-scale video dataset. Concentrating on controllable human-centric generation, pose-guided image-to-video generation [42, 38, 18, 4] adapts a given image to match specific pose sequences. Due to the complex distribution of human video data, they often need to train on extensive high-quality human videos that may be unavailable. To circumvent the need for large-scale human video datasets and extensive training resources, we adopt a one-shot setting. Disco [38], which underwent training on the TikTok dataset [21] offers a strategy for fine-tuning on a single video. In contrast, our method eliminates the need for pre-training on video data, requiring tuning only on a single video of the target individual.
@@ 5.1 Experimental Settings
We train and evaluate our method on test splits of two public datasets TED [34] and TikTok [21]. Due to changes in user privacy and download permissions, we can only download 128 TED test videos from YouTube using download links provided in [34]. Following prior work [38, 42, 4], we use 10 TikTok test videos presented in [38]. We also gather 10 high-quality videos from YouTube, encompassing a variety of scenes such as interviews, movie clips, and talk shows. For each video, we designate 100 frames as test frames and uniformly select $N$ frames from their preceding frames as training frames. We crop and resize all videos to a uniform $512\times 512$ resolution. Please refer to Appendix A.1 for more details and refer to Sec. 6.0.2 for responsibility statements to human subjects.
@@ 6 Conclusions
The video quality of PoseCrafter depends on the ControlNet and latent diffusion models, which struggle with complex poses like crossed fingers, presenting a limitation in our approach. When using poses from others, the body proportions of our generated characters change accordingly. Moreover, large differences between training and inference poses can degrade video quality.
6.0.2 Dataset.
To evaluate the performance of the proposed method on open-domain data, we collected a small-scale dataset consisting of videos with human subjects (see Appendix A.1 for the details) from the Internet. The collected videos were already publicly available online. Our aim is purely academic, and we have not made any offensive edits or alterations to the original content.
6.0.3 Potential Negative Impact.
Appendix
@@ Appendix 0.A Experimental Settings
In this section, we present more experimental details about datasets in Sec. 0.A.1, implementation details in Sec. 0.A.2, and baselines in Sec. 0.A.3.
@@ 0.A.2 Implementation Details
It is worth noting that the aforementioned hyperparameters may not represent the optimal settings, as their detailed tuning was not conducted due to time and computational constraints.
@@ Appendix 0.B Additional Results
In this section, we present more related baselines in Sec. 0.B.1 and more qualitative results in Sec. 0.B.2.
@@ 0.B.1 More Baselines
Tab. 3 and Tab. 4 introduce additional image-to-video baselines, which are marked in gray, for TikTok and TED, respectively. MRAA [34] represents the state-of-the-art GAN-based method and relies on ground truth videos. Consequently, [42] develops an alternative version that requires only DensePose sequences. Moreover, [42] establishes a related baseline, PA+CtrlN-V, integrating existing state-of-the-art image and video generation models. This includes the image-to-image method IP-Adapter [44], which maintains the identity of the reference image, Pose ControlNet [45], controlling the pose in generated videos, and the text-to-video model AnimateDiff [13], ensuring temporal consistency in results. We also report the original quantitative results of Disco and MagicAnimate.

### v3

@@ 1 Introduction
  • We present PoseCrafter, a one-shot pose-guided personalized video generation framework, designed to circumvent the need for large-scale high-quality human video datasets and extensive training resources.
  • PoseCrafter introduces reference-frame selection and insertion to improve sample quality, and further alleviates face and hand degradation through simple latent editing.
@@ 2.1 Controllable Image Generation
Recently, because of the capacity to produce high-fidelity samples, diffusion models are inherently utilized in controllable image generation [51, 16, 10, 1, 53, 27]. Text-prompt-based methods [33, 32, 35] often struggle with precise control, limited by text-image pair quality and complexity of lengthy inputs. Attention control [16, 10] and latent editing [30, 7] can somewhat enhance alignment of text-to-image generation. Personalized methods [34, 19] capture the common concept across a few images and reproduce that concept in generated images. Furthermore, some research [48, 28] introduces adapters that input various conditions like skeleton images to steer semantics and spatial structures. Our work centers on controllable video generation with pre-trained adapters and latent editing.
@@ 2.3 Personalization Video Generation
Text-to-video models [17, 38, 24] create a specific subject video through text description. Image-to-video approaches [2, 12, 29] animate a still image to produce a dynamic variation about it. To harness the vast array of personalized image models in the community, certain techniques [15] transfer these into video generation models by incorporating temporal modeling modules pre-trained on a large-scale video dataset. Concentrating on controllable human-centric generation, some human pose transfer works [50, 37] migrate the motions of a source video to a target subject. However, these methods are inflexible due to using the original video as guidance. More relevant to our work, pose-guided image-to-video generation [45, 41, 20, 6] adapts a given image to match specific pose sequences. Due to the complex distribution of human video data, they often need to train on extensive high-quality human videos that may be unavailable. To circumvent the need for large-scale human video datasets and extensive training resources, we adopt a one-shot setting. Disco [41], which underwent training on the TikTok dataset [23] offers a strategy for fine-tuning on a single video. In contrast, our method eliminates the need for pre-training on video data, requiring tuning only on a single video of the target individual. Similar to our work, Everybody-Dance-Now [5] excludes additional training videos but requires long videos of a target subject (8-17 minutes), while our method merely needs a few video frames (e.g., 8 frames).
@@ 5.1 Experimental Settings
We train and evaluate our method on test splits of two public datasets TED [37] and TikTok [23]. Due to changes in user privacy and download permissions, we can only download 128 TED test videos from YouTube using download links provided in [37]. Following prior work [41, 45, 6], we use 10 TikTok test videos presented in [41]. We also gather 10 high-quality videos from YouTube, encompassing a variety of scenes such as interviews, movie clips, and talk shows. For each video, we designate 100 frames as test frames and uniformly select $N$ frames from their preceding frames as training frames. We crop and resize all videos to a uniform $512\times 512$ resolution. See Appendix A.1 for more details.
@@ 6 Conclusions
The video quality of PoseCrafter depends on the ControlNet and latent diffusion models, which struggle with complex poses like crossed fingers, presenting a limitation in our approach. When using poses from others, the body proportions of our generated characters change accordingly. Moreover, large differences between training and inference poses can degrade video quality. Specific failure cases are presented in Appendix B.4.
Acknowledgements
This work was supported by NSF of China (No. 62076145), Major Innovation & Planning Interdisciplinary Platform for the “Double-First Class" Initiative, Renmin University of China, the Fundamental Research Funds for the Central Universities, the Research Funds of Renmin University of China (No. 22XNKJ13), and the Huawei Group Research Fund. C. Li was also sponsored by Beijing Nova Program (No. 20220484044). The work was partially done at the Engineering Research Center of Next-Generation IntelligentSearch and Recommendation, Ministry of Education.
Supplementary Material
@@ Appendix 0.A Experimental Settings
In this section, we present more experimental details about datasets in Sec. 0.A.1, implementation details in Sec. 0.A.2, and baselines in Sec. 0.A.3. We also detail the inference cost of PoseCrafter in Sec. 0.A.4 and discuss its potential negative impact in Sec. 0.A.5.
@@ 0.A.2 Implementation Details
It is worth noting that the aforementioned hyperparameters may not represent the optimal settings, but we empirically find that they can yield good results as defaults.
@@ 0.A.3 Baselines
0.A.4 Inference Cost
PoseCrafter, with 1.48 billion parameters and around $1.747\times 10^{14}$ FLOPs in total, takes 2.75 GPU minutes to generate 100 frames, utilizing 19.28 GB of memory on a single RTX 3090. In our experiments, using a single RTX 3090 with 24 GB of memory, we successfully generate videos up to a maximum length of 180 frames with good quality. PoseCrafter can generate longer videos with GPUs of larger memory, memory reduction techniques, and long-range video generation strategies.
0.A.5 Potential Negative Impact
@@ Appendix 0.B Additional Results
In this section, we present more related baselines in Sec. 0.B.1, more qualitative results in Sec. 0.B.2, more quantitative results in Sec. 0.B.3, and failure cases of PoseCrafter in Sec. 0.B.4.
@@ 0.B.1 More Baselines
Tab. 3 and Tab. 4 introduce additional image-to-video baselines, which are marked in gray, for TikTok and TED, respectively. TPS [50] and MRAA [37] represents the state-of-the-art GAN-based methods and rely on ground truth videos. Consequently, [45] develops an alternative version that requires only DensePose sequences. Moreover, [45] establishes a related baseline, PA+CtrlN-V, integrating existing state-of-the-art image and video generation models. This includes the image-to-image method IP-Adapter [47], which maintains the identity of the reference image, Pose ControlNet [48], controlling the pose in generated videos, and the text-to-video model AnimateDiff [15], ensuring temporal consistency in results. We also report the original quantitative results of Disco and MagicAnimate.
@@ 0.B.2 More Qualitative Results
0.B.3 More Quantitative Results
We also conduct quantitative experiments on open-domain videos. As shown in Tab. 5, consistent with conclusions on TED and TikTok, PoseCrafter excels in all quantitative metrics than all baselines on our collected open-domain dataset, demonstrating its effectiveness and robustness.
Table 5: Quantitative results on our collected open-domain dataset. NFs represent the number of available frames for the corresponding method. $\dagger$ represents we fine-tune corresponding methods using their tailored strategies for specific subjects. $\ddagger$ indicates corresponding methods directly use ground-truth frames of target poses as input conditions.
Method NFs SSIM $\uparrow$ PSNR $\uparrow$ LPIPS $\downarrow$ FVD $\downarrow$
PIDM [1] 1 0.358 9.36 0.690 1921.43
MRAA [37] 1 0.654 15.26 0.423 1101.64
TPS [50] 1 0.691 17.22 0.317 723.00
DisCo [41] 1 0.589 13.55 0.434 965.47
MagicAnimate [45] 1 0.621 14.84 0.317 518.23
[…]
ControlVideo [52] 8 0.774 20.40 0.174 279.90
PoseCrafter (ours) 8 0.808 21.70 0.143 255.61
0.B.4 Failure Cases
PoseCrafter encounters several limitations in generating videos. Constrained by the capabilities of ControlNet [48] and the latent diffusion model [33], PoseCrafter produces mismatched poses (Fig. 9a) and low-quality results for complex poses (Fig. 9b). A significant discrepancy between training poses and inference poses causes PoseCrafter to yield low-faithfulness videos (Fig. 9c), especially regarding body proportion variation. Moreover, when the training video lacks motion diversity, PoseCrafter is prone to overfitting and struggles with learning temporal consistency.
Figure 9: Failure cases include (a) misalignment of the digital human’s eyes with the derived poses (i.e., closed eyes), (b) poor rendering of the right hand, and (c) changes in the facial structure of the target subject.